A vendor has mangoes that sell for Php 10.00 and Php 12.50 each. How many pieces of each must be mixed together to make 90 pieces of mangoes thet would be sold at Php 11.75 per pieces?
Answer by Maths68(1474)

Number of Kind 1 mangoes = m
Number of Kind 2 mangoes = 90-m
Cost of Kind 1 mango = 10.00
Cost of kind 2 mango = 12.50
Number of mangoes in Mixture = 90
Cost of mango in Mixture = 11.75
(m)(10)+(90-m)(12.5)=11.75*90
10m+1125-12.5m=1057.5
10m-12.5m=1057.5-1125
-2.5m=-67.5
-2.5m/-2.5m=-67.5/-2.5
m=27
Number of Kind 1 mangoes = m=27
Number of Kind 2 mangoes = 90-m=90-27=63